OpenStudy (1645323):

Find the number of real zeros of f(x) = 4x3 + 2 Select one: a. 2 b. 3 c. 4 d. 1

OpenStudy (accessdenied):

You understand that real zeros are solutions to the equation f(x) = 0 such that x is a real number?

OpenStudy (1645323):

yea so how do you solve this?

OpenStudy (accessdenied):

\(4x^3 + 2 = 0\) This is deceptively painless. We can isolate the \(x^3 \) by subtracting the 2 from both sides then dividing by 4, and then we take the cube root of both sides for x.

OpenStudy (accessdenied):

\(x^3 = a \iff x = \sqrt[3]{a}\)

OpenStudy (1645323):

ok so cubedroot(2)?

OpenStudy (accessdenied):

\( 4x^3 + 2 = 0 \) \( 4x^3 = -2 \) Subtract 2 from both sides \( x^3 = \dfrac{-2}{4} \) Divide both ides by 4. Then we can use the cube root. :)

OpenStudy (accessdenied):

It may not seem evident at first why we have a polynomial of degree 3, which we would predict has 3 complex roots, only has this one real value solution. However, we are only using the piece of the cube root that gives us real values (there are two other complex cube roots of a number which we intentionally ignore).

OpenStudy (1645323):

ok so is the answer 2 then?

OpenStudy (accessdenied):

The answer is the quantity of real solutions we found. There could only be one real solution for x^3 = -2/4 <==> x = cuberoot(-2/4)